• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Folts

    Log entry time 14:22:03 on March 16, 2014

    Entry number 386873

    keyword=tech checklist

    A limited checklist that does not include the spectrometers is complete

    Beam line valves upstream of the Compton are open but appear to be disabled

    Beam line valves from Compton to the dump are operational

    The target, spectrometer entrance, spectrometer exit and VDC covers remain installed

    Angle limits for the left HRS is 15 degrees and locked.

    Angle limits for the right HRS is 33 degrees and locked.

    The left and right snout collimators have been removed.

    Both HRS spectrometers are down for troubleshooting of power supplies (inquiries contact Jack Segal)

    There appears to be a solid target in the beam position

    Tech on call is Jessie Butler.